At a spring concert, tickets for adults cost $4.00 and tickets for students cost $2.50. How many of each kind of tickets were purchased if 125 tickets were bought for $413.00?
Answer by mananth(16946)

Adult ticket x numbers
child ticket y numbers

1 x + 1 y = 125 .............1
4 x + 2.5 y = 413 .............2
Eliminate y
multiply (1)by -2.5
Multiply (2) by 1
-2.5 x -2.5 y = -312.5
4 x 2.5 y = 413
Add the two equations
1.5 x = 100.5
/ 1.5
x = 67
plug value of x in (1)
1 x + 1 y = 125
67 + 1 y = 125
1 y = 125 -67
1 y = 58
y = 58
Adult ticket 67 numbers
child ticket 58 numbers
m.ananth@hotmail.ca